what is the length of one leg on an isosceles triangle if the hypotenuse is 10?
Answer by TimothyLamb(4379)

---
h = sqrt( aa + aa )
h = sqrt( 2aa )
hh = 2aa
aa = hh/2
a = sqrt( hh/2 )
a = sqrt( 10*10/2 )
a = 7.071067811865476
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php